If an object has a rest energy of approximately 0.51 MeV, the object is most likely

an electron (m = 9.11 x 10-31 kg).

a proton (m = 1.67 x 10-27 kg).

a carbon atom (m = 1.99 x 10-23 kg).

a molecule of water (m = 2.99 x 10-23 kg).
Answer #1

rest energy = mc^2

where m = mass ,

c = speed of light

0.51*1.6*10^-13 = m*(3*10^8)^2

m = 9.1*10^-31 kg

so this is equal to the mass of electron

so the object is most likely to be an electron